OpenStudy (anonymous):

The vertices of a rectangle are R(5,-5), S(-1,-5) and U(-5,1). After a translation, R' is the point (-11, -11). find the translation rule and coordinates of U'.

OpenStudy (anonymous):

@Opcode ??? Please..

OpenStudy (opcode):

Before translation: R\((5,-5)\) After translation: R\(-11,-11\) \[5 + x = -11\] Solve for \(x\): \[5 + x = -11\] \[x = -16\] ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ \[-5 + y = -11\] Solve for \(y\): \[y = -6\] So the rule is: \[(x,y) \implies (x - 5,~y -6)\] Are you now able to use the rule to find the coordinates of \(U'\)?

OpenStudy (anonymous):

Uhm, I'm not quite sure. Thank you for your help, by the way. I'm super thankful.

OpenStudy (opcode):

\[U' <-5,-6>\] Meaning: \[U'(-5,1) \implies \text{each x and y value goes five towards the left and six down.}\] I do not have graph paper to draw it out but you should be able to understand that, if you don't just ask.

OpenStudy (anonymous):

Okay...I'm still confused on what exactly my answer would be though, I'm so sorry.

OpenStudy (opcode):

The point \(U\) is at \((-5,1)\). According to the rule this point should be translated by: Moving five towards the left and six down. So the final translation point would be: \((-10,-5)\)

OpenStudy (anonymous):

Oh, thank you. Got it.

OpenStudy (opcode):

No problem, glad to help. :-)